OpenStudy (anonymous):

what is 1/3 + blank equals 2

OpenStudy (anonymous):

You just switch around the equation. So 1/3 + x =2 Then - 1/3 on each side x= 5/3 or 1.6666667

OpenStudy (anonymous):

The answer is 5/3 or 1.66667

OpenStudy (anonymous):

5/6

OpenStudy (anonymous):

they are correct